Nordic Journal of Computing

INSTRUCTIONS TO USE NJC STYLE FILES

KIMMO E. E. RAATIKAINEN University of Helsinki, Department of Computer Science

P.O. Box 26 (Teollisuuskatu 23), FIN-00014 University of Helsinki, Finland

               Kimmo.Raatikainen@cs.Helsinki.FI

Abstract. Each paper must include an abstract that summarizes the paper. Recommended length is at most 150 words. The abstract should not contain any references or displayed equations. L^{A}TEX-environment is abstract.

ACM CCS Categories and Subject Descriptors: Each paper should provide the ACM CCS Categories and Subject Descriptors (see e.g. at http://www.acm.org/class/1998/). L^{A}TEX-environment is subject.

Key words: List of keywords (4–6). L^{A}TEX-environment is keywords.

1. Introduction

Nordic Journal of Computing encourages the authors to use L^{A}TEX in preparing the manuscripts with the L^{A}TEX style file njcarticle.cls and the BibTEX style file njcarticle.bst. The figures must be either in the L^{A}TEX picture environment or in Encapsuled Postscript.

This document describes how the L^{A}TEX style file njcarticle.cls and the BibTEX style file njcarticle.bst are to be used. The files can be obtained through anonymous ftp, as follows:

%ftp ftp.cs.Helsinki.FI
Name: anonymous
Password: <your email address>
ftp> cd pub/Nordic_Journal_of_Computing/Style_Files
ftp> get njcarticle.cls
ftp> get njcepsmacros.sty
ftp> get njcarticle.bst
ftp> get njcdoc2e.tex
ftp> get njcdoc.bib
ftp> get njclogo.ps
ftp> quit

The L^{A}TEX-file of an accepted paper, in the final form, should be send through email to njc@cs.Helsinki.FI. If the paper contains Encapsuled Postscript files and/or BibTEX is used send the files separately and an email that explains the files sent.

Received August 20, 1994; revised September 28, 1995; accepted October 11, 1996.

2. General

The L^{A}TEX style file njcarticle.cls does not accept any options. In \documentstyle-command all options given in brackets are simply omitted. The layout of page should not be changed.

Redundant spaces ought to be minimized by careful arrangement of tables and figures. Read your .log file carefully. Remember there should be no Overfull \hbox (as here) and certainly no visible one (more than 1pt). If necessary, reword the text. The preamble command \draft can be used to produce a visible overfullrulebox in the margin.

The style file provides macros to create running heads. The \authorhead contains the authors’ names as “Fistname Surname”, “F. Surname1, F. Surname2”, and “F. Surname1, F. Surname2, F. Surname3” for up to three authors, and “Surname1 et al.” for four or more authors. The \titlehead contains a short form of the title, not more than 30 characters.

The argument of title-command must be written in capitals. When the paper has more than one author, the authors in the argument of author-command are separated by usual \and-command or by \AND-command that inserts vertical glue between the blocks of authors’ names, affiliations, etc.

Neither footnotes nor appendices should be used. However, if appendices are really necessary, their place is after Acknowledgements and before References.

3. Sectioning, numbering, etc.

3.1 Sectioning

The following three L^{A}TEX-sectioning commands are available:

\section, \subsection, and \subsubsection.

3.2 Numbering

The numbering of displayed equations, theorems, figures, tables, and other “numbered” environments follows one of two styles: either consecutive in each section, or consecutive through the whole paper (default). If you prefer the first style, the style file provides \twolevelnumbering-command. The argument of the command is a list of environments that are numbered consecutively in each section, e.g.

    \twolevelnumbering{figure,equation,theorem}

Tables should be referred to “Table I”. Equations and figures should be referred to in abbreviated forms: “Eq. (4.1)” and “Fig. 5.1”. Use macros \eqnref{eqnlabel}, \figref{fig-label}, and \tableref{tab-label} to generate the references.

3.3 Lists of items

The depths of L^{A}TEX-environments itemize and enumerate are restricted to two.

3.4 Spacing before and after environments

Extra space is added at the top of list if the input file has a blank line before any listmaking environment. The vertical space after the environment is the same as the one preceding it. The list-making environments are: quote, quotation, verse,

itemize, enumerate, description, center, flushleft, and flushright, as

well as the theorem-like environments.

3.5 Tabbing environment

The style file defines \tabbingstretch to specify the strut to be used in the tabbing environment. The functionality of \tabbingstretch is the same as

that of \arraystretch in array and tabular environments.

4. Theorem-like environments

The style file defines the following theorem-like environments:

         theorem    proposition  claim    fact
         lemma      definition   problem  remark
         corollary  conjecture   example  observation

These are “numbered” environments. The style file also defines the corresponding “unnumbered” environments: theorem*, ..., observation*. The proofs can be typed within the environment proof.

For example, the L^{A}TEX source

\begin{theorem}[Chebychev's Inequality]
If $X$ is any random variable, then
\begin{equation}
\Pr[|X|\ge a] \le \mbox{\rm E}(X^2)/a^2\;.
\label{eqn:example}
\end{equation}
\end{theorem}

produces

Theorem 1. (Chebychev’s Inequality)If X is any random variable, then

Pr[|X| ≥ a] ≤ E(X^{2})/a^{2} .  (4.1)

Proof. If F(x) denotes the distribution function of the random variable X, then

Pr[|X| ≥ a] = ∫_{|x|≥a} d F(x) .

Since in the region of integration |x|/a ≥ 1, it follows that

∫_{|x|≥a} d F(x) ≤ \frac{1}{a^{2}} ∫_{|x|≥a}x^{2} d F(x) .

By extending the integration to all values of x, we merely strengthen the inequality:

∫_{|x|≥a} d F(x) ≤ \frac{1}{a^{2}} ∫_{|x|≥a}x^{2} d F(x) ≤ \frac{1}{a^{2}} ∫ x^{2} d F(x) = E(X^{2})/a^{2} .

When a proof ends with a displayed equation as above, the box “ ” should be at the right end of the formula rather than at the beginning of the next line. In this case, the box must be inserted through \qed at the end of the equation, and a blank line must be left between the closing of the equation and the \end{proof}. (See the source of these instructions.)

5. Figures and tables

Figures and tables are to be inserted in the text nearest their first reference. They should be arranged so as not to cause an excessive amount of blank space on the remainder of the page.

The captions are centered below the figures and above the tables. If a table needs to extend over to a second page, the continuation of the table should have a caption: “Table II (cont.)”. Macro \continued generates this caption. Example:

     \begin{table}
     \continued   % Instead of \caption
        ...
     \end{table}

Fig. 5.1 is an example of inserting an Encapsuled Postscript file. The style file provides the command \usenjcepsmacros to read the option file njcepsmacros.sty. The Encapsuled Postscript file is inserted using the command:

        \makeepsbox[pos]{filename}[scaling]

The optional argument pos tells how the box is placed on the page:

c centered (default)
l leftmargin justified
r rightmargin justified

The optional argument scaling tells how the eps-file is to be scaled. The possible values of the argument are:

 scale=<factor> scaling factor = <factor>/1000
 width=<dimen> scaling factor = <dimen>/<natural width>
 height=<dimen> scaling factor = <dimen>/<natural height>

Since the TEX-command \special is used, all device drivers may not understand the format. If your driver has problems, you can use \epsdraft-command in the preamble. This command changes the \makeepsbox-command to omit the \special-command. In this case the only visible effect is reservation of space.

Observe that an Encapsuled Postscript file must begin with “%!” and must contain a commentline

   %%BoundingBox: llx lly urx ury

Fig. 5.1: An example.

If the BoundingBox-comment is not found, the file is omitted.

Remember that previously published material must be accompanied by written permission from the author and publisher!

6. References

The preferred style of referring to the bibliography is: “Raatikainen [1993] proposed to use the Bonferroni inequality [see Kleijnen 1987, pp. 41–43; Law and Kelton 1991, Ch. 9.4].”

Note that the years are in brackets in the running text, but without brackets if the reference itself is in brackets. For three or more authors, use et al. Several papers in the same year are distinguished as “Raatikainen 1994a, Raatikainen 1994b”. When an article has more than two authors, the citation should be Kojo et al. [1996].

The second possible style is to use numbered references and the standard L^{A}TEX-

environment thebibliography.

The style file provides the command \njccitationstyle to activate the preferred citation style. The following macros are defined to simplify the preferred style:

macro  procudes

\cite{bibreflabel}  “Author [Year]”
\cite[note]{bibreflabel}  “Author [Year, note]”
\shortcite{bibreflabel}  “Author Year”
\shortcite[note]{bibreflabel}  “Author Year, note”
\citeauthor{bibreflabel}  “Author”
\citeyear{bibreflabel}  “Year”
\bracketcite{bibreflabel}  “[Author Year]”
\bracketcite[note]{bibreflabel}  “[Author Year, note]”
\citemul{bibreflabel1,bibreflabel2} “Author [Year1, Year2]”

The macros above assume that your bibliography items are written as

\bibitem[{Author}{Year}]{bibreflabel}
\bibitem[{Author}{Year1}]{bibreflabel1}
\bibitem[{Author}{Year2}]{bibreflabel2}

The BibTEX style file njcarticle.bst generates the bibliography items in the format above.

Observe that the \citemul-command assumes that the authors of the references are one person (or same persons): Raatikainen [1993, 1994a, 1994b].

7. Use of the style files

Nordic Journal of Computing provides the following files:

(1) njcarticle.cls contains the L^{A}TEX code for producing camera-ready output accepted by Nordic Journal of Computing.

(2) njcepsmacros.sty contains the L^{A}TEX code for including Encapsuled Postscript in the camera-ready output.

(3) njcarticle.bst contains the BibTEX code for producing a reference list suitable for Nordic Journal of Computing.

(4) njcdoc2e.tex contains the L^{A}TEX-source of these instructions.

(5) njcdoc.bib contains the BibTEX-source of these instructions.

(6) njclogo.ps contains the Encapsuled Postscript file used in these instructions.

To format the document, type:

 latex njcdoc2e
 bibtex njcdoc2e
 latex njcdoc2e
 latex njcdoc2e

8. Disclaimer

The macros are not guaranteed to be free of errors. Any bugs, inconsistencies, suggestions, and other comments should be reported to Nordic Journal of Computing

by email to njc@cs.Helsinki.FI.

Acknowledgements

This section comes before the References and is unnumbered. L^{A}TEX-environment

is acknowledgements.

References

Kleijnen, J. P. C. 1987. Statistical Tools for Simulation Practitioners. Marcel Dekker, Inc., New

York. Kojo, M., Raatikainen, K., and Alanko, T. 1996. Connecting Mobile Workstations to the Internet

over a Digital Cellular Telephone Network. In Mobile Computing, Imieliński, T. and Korth, H. F., Editors. Kluwer, Boston, Mass., 201–55.

Law, A. M. and Kelton, W. D. 1991. Simulation Modeling & Analysis, 2nd edition. McGraw-Hill,

New York. Raatikainen, K. E. E. 1993. Sequential Procedure for Simultaneous Estimation of Several Means.

ACM Transactions on Modeling and Computer Simulation 3, 2 (Apr.), 108–133. Raatikainen, K. E. E. 1994a. Database Access in Intelligent Networks. In Proceedings of IFIP

TC6 Workshop on Intelligent Networks. Lappeenranta University of Technology, Lappeenranta, Finland, 163–183.

Raatikainen, K. E. E. 1994b. Information Aspects of Services and Service Features in Intelligent Network Capability Set 1. Report C-1994-45, University of Helsinki, Dept. of Computer Science, Helsinki, Finland.